The probability of a light bulb being defective is 0.02. If Home Depot buys 100 bulbs, find the probability that they get at
least one that is defective.

Answer by rothauserc(4718)

this uses binomial probability
p = .02, q = .98, n = 100
**********************************************
binomial probability formula is
Probability ( X = k ) = (combination of n things taken k at a time) * p^k * q^(n-k)
**********************************************
Probability ( X > or = 1 ) = 1 - P( X = 0 )
P( X = 0 ) = combination of 100 light bulbs taken 0 at a time) * (.02)^0 * (.98)^100 = (( 100! / (0! * (100 - 0)!)) * 1 * 0.132619556
P( X = 0 ) = 1 * 0.132619556 approx 0.13
Probability ( X > or = 1 ) = 1 - 0.13 = 0.87
therefore,
probability that they get at least one that is defective is 0.87